### v1

@@ 3 Experiment
3.1 Experimental Setup
Evaluation Models.
In the evaluation of LMMs, we selected over 50 models that exhibited strong performance on the VLM leaderboard. We mainly focus on open-source models, which include Qwen2-VL [38], InternVL2 [11], LLaVA-OV [25], Ovis [31], etc. Additionally, we evaluated three closed-source API models: GPT-4o, Qwen-VL-Max, and Qwen-VL-Plus. To process the image-prompt generated by LMMs and create new images, we selected four text-to-image generation models: FLUX.1-dev, Stable Diffusion 3.5, Kolors and Lumina. During the final evaluation phase, the Unicom image representation model was utilized to extract image features. Further details about our baseline models are provided in the Appendix C.2.
Implementation Details.
During the image-prompt generation process in LMMs, we utilize the default settings of VLMEvalKit [14], modifying only the predefined query prompt and the image to meet the task-specific requirements. Subsequently, we employ four distinct text-to-image models to generate new images and extract features using an image representation model. We then compute both the SIM-Score and FID-Score for evaluation. Our analysis, detailed in Appendix C.2, reveals a high degree of consistency among the results produced by the four generative models. Therefore, unless otherwise stated, the reported results are obtained using the FLUX.1-dev text-to-image model.
@@ 3.2 Main Results
Overall Performance.
As shown in Table 1 and 2, the SIM-Score of the most advanced LMMs on MMGenBench-Test is below $0.600$. Specifically, GPT-4o achieved a score of $0.566$, which is lower than the $0.599$ obtained by the open-source model InternVL2-76B. Notably, there is no clear correlation between the model size and performance across different series of models. This indicates the importance of training data and the training process in developing LMMs with strong capabilities of image understanding and description. Furthermore, it is observed that models excelling on existing benchmarks do not necessarily perform well on MMGenBench-Test, such as LLaVA-OV, which only scores $0.494$. In the MMGenBench-Domain dataset, the SIM-Score results align closely with those of the MMGenBench-Test. However, there is a significant difference in FID-Score because MMGenBench-Domain includes 10,000 images, thereby improving the accuracy of its FID-Score measurement. Therefore, we propose using SIM-Score as the primary metric. Detailed results and comprehensive analysis are provided in the following sections as well as in Appendix C.3.
@@
Image Patterns Revealing Model’s Strengths and Limitations.
Fig. 5 shows the accuracy of the best-performing LMMs on the MMGenBench dataset. We observed that LMMs exhibit superior performance on image patterns such as “Artistic”, “Surreal”, “Symbol”, “Color”, etc. Conversely, their performance declines on patterns such as “Contextual”, “Orientation”, “Count”, “Motion”, etc. This discrepancy suggests that LMMs are proficient in tasks requiring coarse-grained image understanding and description. In contrast, their abilities for fine-grained understanding and description are more challenging, as these tasks require the understanding of the complex contextual relationships within images and the precise linguistic description.
Impact of Parameter Scales, Training Methods, and Data Variations.
Fig.6(a) shows the performance results of the same series of models with varying parameter sizes, while Fig.6(b) presents the performance results of the same model and parameters under different training protocols and datasets. We can see that as the model parameters increase, the performance of Qwen2-VL improves from $0.487$ to $0.553$, and the results of InternVL2 increased from $0.476$ to $0.599$. Additionally, by using improved training data and protocols, Ovis1.6 outperforms Ovis1.5 by $0.06$. Given that our evaluation is based on a single input image, LLaVA-OV-SI demonstrates superior performance to LLaVA-OV.
@@ 3.3 Analysis
Failure in Following Instructions.
As shown in Fig. 7, we have selected various LMMs to demonstrate a portion of the qualitative examples. It is observed that certain LMMs fail to strictly follow instructions when generating the image-prompt. Specifically, some LMMs prepend a prefix to the image-prompt (e.g., RBDash-72B [3] in Fig. 7), while others append explanatory content following the image-prompt (e.g., MiniCPM-Llama3-V2.5 [44] in Fig. 7). Additionally, it is also evident that the ability to follow instructions does not appear to be correlated with the parameter size of different series LMMs. For example, both Qwen2-VL-2B and InternVL2-2B accurately generate the image-prompt in accordance with the given instructions. We argue that an effective LMM, especially when undergoing instruction tuning, should be capable of accurately following instructions and successfully completing the task.
Inability in Generating Detailed Image-prompt.
Most of the image-text pairs used in the training of existing LMMs have relatively short image captions. Consequently, numerous models exhibit suboptimal performance on the MMGenBench Benchmark, which requires the model to understand the image and provide a detailed description. As shown in Fig. 7, LLaVA-OV-72B [25], despite its extensive 72B parameters, produces relatively short image-prompt, which hinders its ability to excel on the MMGenBench Benchmark. Additionally, we observed that for LMMs in the same series, smaller LMMs tend to generate shorter image-prompt, as evidenced in the Qwen2-VL and InternVL2 series. Based on these findings, we hypothesize that incorporating image-description pairs during the training phase of LMMs could achieve superior results.
Model Overfitting.
Existing LMMs frequently undergo task-specific training to meet established benchmarks. This specialized training can lead to overfitting, which may compromise the basic image understanding and description capabilities of LMMs. For instance, the xGen-MM [43] results in Fig. 7 demonstrate an overfitting towards the notion of “safety”, despite the given image not presenting any safety risk. This observation suggests that in adapting models to various tasks, it is crucial to maintain the basic image understanding and description capabilities of LMMs.
@@ 4 Conclusion
In this study, we present a straightforward automated evaluation pipeline to comprehensively evaluate the ability of LMMs in image understanding and description. Based on this pipeline, we introduce MMGenBench-Test and MMGenBench-Domain to evaluate LMM performance across different image patterns and domain-specific images. By evaluating over 50 widely-used LMMs, we demonstrate the reliability and effectiveness in both the pipeline and benchmarks. In conclusion, our method provides a more fundamental evaluation of existing LMMs, identifies specific gaps in their image understanding and description capabilities, and establishes a robust foundation for further research and model improvement in these areas. All our data and code are open source.
@@ A.1 Automatic Benchmarks
The rapid advancements in LLMs have prompted the development of diverse benchmarks to evaluate their performance. Early efforts, such as LMExamQA  [8], introduced the ”Language-Model-as-an-Examiner” framework to provide a scalable and comprehensive evaluation for LLMs. However, there remains a significant gap in evaluating the reasoning capabilities of LLMs, especially in dynamic contexts. To address this deficiency, recent frameworks like DYVAL  [51] and DYVAL2  [52] have focused on dynamic reasoning tasks. DYVAL concentrates on reasoning abilities, while DYVAL2 extends the evaluation to a broader psychometric approach, thereby providing deeper insights into cognitive evaluations. Additionally, AutoBencher  [26] has automated the generation of challenging and novel datasets, specifically designed for LLM evaluation. Platforms such as UNIGEN  [41] and Task Me Anything  [49] aim to further enhance evaluation by developing domain-specific benchmarks tailored to the unique capabilities of LLMs.
@@ A.2 LMM Benchmarks
The emergence of LMMs has highlighted the insufficiency of traditional benchmarks, which primarily focus on isolated task performance and fail to capture the full capabilities exhibited by multimodal models. Early multimodal benchmarks, such as those introduced by [27, 19, 48], established essential foundations but lacked the granularity required for robust evaluation of the model’s understanding and reasoning abilities. Recent studies [30, 10, 47] emphasize the critical need for comprehensive, fine-grained benchmarks that can more accurately evaluate the complex capabilities of LMMs. However, existing benchmarks, such as LVLM-eHub [42] and LAMM [45], still rely on classical datasets that do not fully reflect the current state of the field, and neglect issues such as data leakage during model training. Addressing these concerns, MMStar [10] introduces an innovative approach that eliminates visual content dependencies and mitigates data leakage risks, thereby providing a more reliable and secure evaluation framework. Furthermore, the development of automated benchmarking systems, such as AUTOBENCH-V[9], enhances the scalability and objectivity of evaluations, by automating the benchmarking process and minimizing human biases. Despite these significant advancements, further refinement is needed to create truly comprehensive and dynamic benchmarks that can capture the evolving capabilities of LMMs.
@@ B.1 Details of MMGenBench-Test Data
Thirteen Image Patterns.
As illustrated in Fig. 10, summarization is performed using GPT-4 Turbo, followed by human verification. We identify $13$ distinct image patterns, each with a comprehensive explanation. It should be noted that a single image can exhibit multiple patterns concurrently.
Patterns Extracted by GPT-4o.
Fig. 11 shows the image patterns extracted by GPT-4o, organized in descending order. While the total number of extracted patterns is 1868, only the most frequently occurring patterns are presented due to space constraints.
@@ B.3 Details of Construction Prompts
To construct the MMGenBench-Test, a series of prompts have been tailored to the specific requirements of our task meticulously. Fig. 12, 13, and 14 include the three sequential stages in the pipeline (see Fig. 2) used to develop MMGenBench-Test.
@@ C.1 Evaluation Pipeline Prompt
We meticulously crafted a prompt to guide LMMs in generating image-prompt for subsequent evaluation. As shown in Fig. 15, the prompt encompasses five key perspectives: “role”, “definition”, “task”, “key points”, and “output”. To ensure the objectivity and consistency of the evaluation, we employed this singular prompt, as it is sufficiently clear to enable the model to effectively perform the image-prompt generation task.
@@ C.2 Experimental Setup
We employed over 50 popular LMMs, as detailed in Tables 3 and 4, which are available on the OpenVLM Leaderboard [13]. For open-source LMMs, we utilized the original configuration of VLMEvalKit [14] to perform inference of image-prompt on $8$ NVIDIA-H20 GPUs. During the image generation and metric calculation stages, we still use $8$ NVIDIA-H20 GPUs for multi-process inference to efficiently complete the task.
We analyzed the performance metrics of four text-to-image models on the MMGenBench-Test dataset, as illustrated in Fig. 9. The results indicate that the four text-to-image models exhibited consistency in both SIM-Score and FID-Score. Consequently, we report the results of FLUX.1-dev as the final results in the paper.
@@ Appendix D Qualitative Example
Qualitative analyses for more than 50 popular LMMs are illustrated in Fig. 16-20.
@@ Appendix E Limitations and Future Work
Our method leverages the powerful capabilities of text-to-image and image representation models. Despite our rigorous efforts to minimize errors, the evaluation pipeline may not achieve complete accuracy for tasks, that present significant challenges to certain text-to-image and image representation models, such as those involving table images. Our research has the potential to inform and refine future methodologies for modifying target tasks and evaluation techniques, specifically addressing challenges faced by current text-to-image and image representation models. For example, in the context of table images, LaTeX could be utilized to render content and facilitate pixel-level image matching.
@@
SIM FID SIM FID
Qwen2-VL-72B-Instruct 0.553 1.357 0.545 0.710
qwen-vl-max-0809 0.552 1.363 - -
InternVL2-76B 0.599 1.264 0.599 0.632
GPT4o_HIGH 0.566 1.306 - -
InternVL2-40B 0.566 1.350 0.566 0.696
Ovis1.6-Gemma2-9B 0.582 1.316 0.579 0.667
llava_onevision_qwen2_72b_ov 0.494 1.561 0.491 0.872
llava_onevision_qwen2_72b_si 0.512 1.512 0.514 0.813
Qwen2-VL-7B-Instruct 0.532 1.437 0.524 0.775
InternVL2-26B 0.576 1.320 0.577 0.671
qwen-vl-plus 0.475 1.586 - -
MiniCPM-V-2_6 0.548 1.386 0.545 0.710
InternVL2-8B 0.547 1.403 0.548 0.701
Ovis1.5-Gemma2-9B 0.521 1.487 0.524 0.808
RBDash_72b 0.525 1.413 0.527 0.740
MMAlaya2 0.537 1.423 0.541 0.747
Ovis1.5-Llama3-8B 0.526 1.466 0.527 0.795
OmChat 0.439 1.761 0.447 1.028
InternVL-Chat-V1-5 0.547 1.393 0.547 0.729
llava_onevision_qwen2_7b_si 0.492 1.554 0.497 0.834
XComposer2d5 0.460 1.646 0.458 0.955
Pixtral-12B 0.552 1.372 0.561 0.686
InternVL2-4B 0.556 1.345 0.556 0.689
llava_onevision_qwen2_7b_ov 0.488 1.590 0.490 0.861
glm-4v-9b 0.511 1.463 0.527 0.730
molmo-7B-D-0924 0.519 1.490 0.523 0.813
XComposer2_4KHD 0.433 1.781 0.435 1.075
MiniCPM-Llama3-V-2_5 0.526 1.432 0.530 0.767
VILA1.5-40b 0.483 1.656 0.486 0.958
cambrian_34b 0.546 1.377 0.549 0.701
WeMM 0.488 1.565 0.479 0.888
Llama-3.2-11B-Vision-Instruct 0.526 1.428 0.525 0.768
Qwen2-VL-2B-Instruct 0.487 1.549 0.501 0.806
XComposer2 0.248 2.890 0.247 2.130
cogvlm2-llama3-chat-19B 0.465 1.523 0.464 0.825
Idefics3-8B-Llama3 0.460 1.718 0.458 1.004
Mini-InternVL-Chat-4B-V1-5 0.527 1.433 0.525 0.765
XinYuan-VL-2B-Instruct 0.395 1.951 0.392 1.240
llava_next_yi_34b 0.476 1.636 0.483 0.883
InternVL2-2B 0.476 1.563 0.483 0.848
Phi-3-Vision 0.457 1.678 0.461 0.960
cambrian_13b 0.509 1.526 0.506 0.852
Phi-3.5-Vision 0.479 1.593 0.478 0.892
idefics2_8b 0.449 1.761 0.453 1.015
cambrian_8b 0.512 1.498 0.515 0.810
minimonkey 0.450 1.673 0.455 0.939
Llama-3-MixSenseV1_1 0.474 1.615 0.477 0.922
llava_next_qwen_32b 0.507 1.501 0.515 0.796
xgen-mm-phi3-interleave-r-v1.5 0.529 1.454 0.529 0.761
Eagle-X5-13B 0.434 1.812 0.446 1.050
xgen-mm-phi3-dpo-r-v1.5 0.414 1.671 0.412 0.940
Mini-InternVL-Chat-2B-V1-5 0.428 1.709 0.440 0.973
llava_next_llama3 0.476 1.601 0.486 0.881
VILA1.5-13b 0.485 1.695 0.480 1.006
Mantis-8B-siglip-llama3 0.411 1.833 0.409 1.116

### v2

@@ 4 Experiment
4.1 Experimental Setup
Evaluation Models. In the evaluation of LMMs, we selected over 50 models that exhibited strong performance on the VLM leaderboard [15]. We mainly focus on open-source models, which include Qwen2-VL [65], InternVL2 [10], LLaVA-OV [37], Ovis [50], etc. Additionally, we evaluated three closed-source API models: GPT-4o [53], Qwen-VL-Max [6], and Qwen-VL-Plus [6]. To process the textual descriptions generated by LMMs and create auxiliary images, we selected four text-to-image generation models: FLUX.1-dev [34], Stable Diffusion 3.5 [16], Kolors [62] and Lumina [20]. During the final evaluation phase, the Unicom [5] image representation model was utilized to extract image features. Further details about our baseline models are provided in the Appendix C.2.
Implementation Details. During the textual descriptions generation process in LMMs, we utilize the default settings of VLMEvalKit [15], modifying only the predefined query prompt and the image to meet the task-specific requirements. Subsequently, we employ four distinct text-to-image models to generate auxiliary images and extract features using an image representation model. We then compute both the SIM-Score and FID-Score for evaluation. We analyzed the performance metrics of four text-to-image models on the MMGenBench-Test dataset, as illustrated in Fig. 6. The results indicate that the four text-to-image models exhibited consistency in both SIM-Score and FID-Score. Therefore, unless otherwise stated, the reported results are obtained using the FLUX.1-dev text-to-image model. Please refer to Appendix C.2 for more detailed results.
@@ 4.2 Main Results
Overall Performance. As shown in Table 1 and 2, the SIM-Score of the most advanced LMMs on MMGenBench-Test is below $0.600$. Specifically, GPT-4o achieved a score of $0.566$, which is lower than the $0.599$ obtained by the open-source model InternVL2-76B. Notably, there is no clear correlation between the model size and performance across different series of models. This indicates the importance of training data and the training process in developing LMMs with strong capabilities of image understanding and description. Furthermore, it is observed that models excelling on existing benchmarks do not necessarily perform well on MMGenBench-Test, such as LLaVA-OV, which only scores $0.494$. In the MMGenBench-Domain dataset, the SIM-Score results align closely with those of the MMGenBench-Test. However, there is a significant difference in FID-Score because MMGenBench-Domain includes 10,000 images, thereby improving the accuracy of its FID-Score measurement. Therefore, we propose using SIM-Score as the primary metric. Detailed results and comprehensive analysis are provided in the following sections as well as in Appendix C.3.
@@
Image Patterns Revealing Model’s Strengths and Limitations. Fig. 7 shows the accuracy of the best-performing LMMs on the MMGenBench dataset. We observed that LMMs exhibit superior performance on image patterns such as “Artistic”, “Surreal”, “Symbol”, “Color”, etc. Conversely, their performance declines on patterns such as “Contextual”, “Orientation”, “Count”, “Motion”, etc. This discrepancy suggests that LMMs are proficient in tasks requiring coarse-grained image understanding and description. In contrast, their abilities for fine-grained understanding and description are more challenging, as these tasks require the understanding of the complex contextual relationships within images and the precise linguistic description.
Impact of Parameter Scales, Training Methods, and Data Variations. Fig.8(a) shows the performance results of the same series of models with varying parameter sizes, while Fig.8(b) presents the performance results of the same model and parameters under different training protocols and datasets. We can see that as the model parameters increase, the performance of Qwen2-VL improves from $0.487$ to $0.553$, and the results of InternVL2 increased from $0.476$ to $0.599$. Additionally, by using improved training data and protocols, Ovis1.6 outperforms Ovis1.5 by $0.06$. Given that our evaluation is based on a single input image, LLaVA-OV-SI demonstrates superior performance to LLaVA-OV.
@@ 4.3 Analysis
Failure in Following Instructions. As shown in Fig. 9, we have selected various LMMs to demonstrate a portion of the qualitative examples. It is observed that certain LMMs fail to strictly follow instructions when generating the textual descriptions. Specifically, some LMMs prepend a prefix to the textual descriptions (e.g., RBDash-72B [57] in Fig. 9), while others append explanatory content following the textual descriptions (e.g., MiniCPM-Llama3-V2.5 [75] in Fig. 9). Additionally, it is also evident that the ability to follow instructions does not appear to be correlated with the parameter size of different series LMMs. For example, both Qwen2-VL-2B and InternVL2-2B accurately generate the textual descriptions following the given instructions. We argue that an effective LMM, especially when undergoing instruction tuning, should be capable of accurately following instructions and successfully completing the task.
Inability in Generating Detailed Textual Descriptions. Most of the image-text pairs used in the training of existing LMMs have relatively short image captions. Consequently, numerous models exhibit suboptimal performance on the MMGenBench Benchmark, which requires the model to understand the image and provide a detailed description. As shown in Fig. 9, LLaVA-OV-72B [37], despite its extensive 72B parameters, produces relatively short textual descriptions, which hinders its ability to excel on the MMGenBench Benchmark. Additionally, we observed that for LMMs in the same series, smaller LMMs tend to generate shorter textual descriptions, as evidenced in the Qwen2-VL and InternVL2 series. Based on these findings, we hypothesize that incorporating image-description pairs during the training phase of LMMs could achieve superior results.
Model Overfitting. Existing LMMs frequently undergo task-specific training to meet established benchmarks. This specialized training can lead to overfitting, which may compromise the basic image understanding and description capabilities of LMMs. For instance, the xGen-MM [73] results in Fig. 9 demonstrate an overfitting towards the notion of “safety”, despite the given image not presenting any safety risk. This observation suggests that in adapting models to various tasks, it is crucial to maintain the basic image understanding and description capabilities of LMMs.
@@ 5 Conclusion
In this paper, we present a straightforward and fully automated evaluation pipeline to comprehensively evaluate the ability of LMMs in image understanding and description. Based on MMGenBench-pipeline, we introduce MMGenBench-Test and MMGenBench-Domain to evaluate LMM performance across different image patterns and domain-specific images. By evaluating over 50 widely used LMMs, we demonstrate the reliability and effectiveness of both the pipeline and benchmarks. In conclusion, our method provides a more fundamental evaluation of existing LMMs, identifies specific gaps in their image understanding and description capabilities, and establishes a robust foundation for further research and model improvement in these areas. All code and data will be released soon.
@@ A.1 Automatic Benchmarks
The rapid advancements in LLMs have prompted the development of diverse benchmarks to evaluate their performance. Early efforts, such as LMExamQA  [7], introduced the ”Language-Model-as-an-Examiner” framework to provide a scalable and comprehensive evaluation for LLMs. However, there remains a significant gap in the evaluation of the reasoning capabilities of LLMs, especially in dynamic contexts. To address this deficiency, recent frameworks like DYVAL  [83] and DYVAL2  [84] have focused on dynamic reasoning tasks. DYVAL concentrates on reasoning abilities, while DYVAL2 extends the evaluation to a broader psychometric approach, thereby providing deeper insights into cognitive evaluations. Additionally, AutoBencher  [38] has automated the generation of challenging and novel datasets, specifically designed for the evaluation of LLM. Platforms such as UNIGEN  [69] and Task Me Anything  [80] aim to further enhance evaluation by developing domain-specific benchmarks tailored to the unique capabilities of LLMs.
@@ A.2 LMM Benchmarks
The emergence of LMMs has highlighted the insufficiency of traditional benchmarks, which primarily focus on isolated task performance and fail to capture the full capabilities exhibited by multimodal models. Early multimodal benchmarks, such as those introduced by [41, 23, 79], established essential foundations but lacked the granularity required for robust evaluation of the model’s understanding and reasoning abilities. Recent studies [46, 9, 78] emphasize the critical need for comprehensive, fine-grained benchmarks that can more accurately evaluate the complex capabilities of LMMs. However, existing benchmarks, such as LVLM-eHub [72] and LAMM [76], still rely on classical datasets that do not fully reflect the current state of the field, and neglect issues such as data leakage during model training. Addressing these concerns, MMStar [9] introduces an innovative approach that eliminates visual content dependencies and mitigates data leakage risks, thereby providing a more reliable and secure evaluation framework. Furthermore, the development of automated benchmarking systems, such as AUTOBENCH-V[8], enhances the scalability and objectivity of the evaluations, by automating the benchmarking process and minimizing human bias. Despite these significant advancements, further refinement is needed to create truly comprehensive and dynamic benchmarks that can capture the evolving capabilities of LMMs.
@@ B.1 Details of MMGenBench-Test Data
Thirteen Image Patterns. As illustrated in Fig. 11, summarization is performed using GPT-4 Turbo, followed by human verification. We identify $13$ distinct image patterns, each with a comprehensive explanation. It should be noted that a single image can exhibit multiple patterns concurrently.
Patterns Extracted by GPT-4o. Fig. 12 shows the image patterns extracted by GPT-4o, organized in descending order. While the total number of extracted patterns is 1868, only the most frequently occurring patterns are presented due to space constraints.
@@ B.3 Details of Construction Prompts
To construct the MMGenBench-Test, a series of prompts have been tailored to the specific requirements of our task meticulously. Fig. 13, 14, and 15 include the three sequential stages in the pipeline (see Fig. 3) used to develop MMGenBench-Test.
@@ C.1 Evaluation Pipeline Prompt
We meticulously crafted a prompt to guide LMMs in generating textual descriptions of images for subsequent evaluation. As shown in Fig. 16, the prompt encompasses five key perspectives: “role”, “definition”, “task”, “key points”, and “output”. To ensure the objectivity and consistency of the evaluation, we employed this singular prompt, as it is sufficiently clear to enable the model to effectively perform the textual description generation task.
@@ C.2 Experimental Setup
We employed over 50 popular LMMs, as detailed in Tables 3 and 4, which are available on the OpenVLM Leaderboard [13]. For open-source LMMs, we utilized the original configuration of VLMEvalKit [15] to perform inference of textual descriptions of images on $8$ NVIDIA-H20 GPUs. During the image generation and metric calculation stages, we still use $8$ NVIDIA-H20 GPUs for multi-process inference to efficiently complete the task.
@@ Appendix D Qualitative Example
Qualitative analyses for more than 50 popular LMMs are illustrated in Fig. 17-21.
@@ Appendix E Limitations and Future Work
Our MMGenBench-pipeline leverages the powerful capabilities of text-to-image and image representation models. Despite our rigorous efforts to minimize errors, the evaluation pipeline may not achieve complete accuracy for tasks, which presents significant challenges to certain text-to-image and image representation models, such as those involving table images. Our research has the potential to inform and refine future methodologies for modifying target tasks and evaluation techniques, specifically addressing challenges faced by current text-to-image and image representation models. For example, in the context of table images, LaTeX could be utilized to render content and facilitate pixel-level image matching.
In addition, the adoption of the text-to-image model and the image representation model in our MMGenBench-pipeline has led to an increase in evaluation time. Despite this, it is important to note that the MMGenBench-pipeline remains more cost-effective compared to manually constructed benchmarks. Moreover, our MMGenBench-pipeline is fully automated and eliminates the need for human intervention in the evaluation of LMM performance across various domains, making it inherently more scalable than previous benchmarks. This robust scalability enables our pipeline to seamlessly adapt to more powerful models and updated datasets. As the advancements in text-to-image models continue, the time efficiency challenges we currently encounter are expected to improve. For instance, the flux-schnell requires only four steps to generate images, which can significantly accelerate our evaluation process.
@@
[…]
Qwen2-VL-72B-Instruct [65] 0.553 1.357 0.545 0.710
qwen-vl-max-0809 [6] 0.552 1.363 - -
InternVL2-76B [10] 0.599 1.264 0.599 0.632
GPT4o_HIGH [53] 0.566 1.306 - -
InternVL2-40B [10] 0.566 1.350 0.566 0.696
Ovis1.6-Gemma2-9B [50] 0.582 1.316 0.579 0.667
llava_onevision_qwen2_72b_ov [37] 0.494 1.561 0.491 0.872
llava_onevision_qwen2_72b_si [37] 0.512 1.512 0.514 0.813
Qwen2-VL-7B-Instruct [65] 0.532 1.437 0.524 0.775
InternVL2-26B [10] 0.576 1.320 0.577 0.671
qwen-vl-plus [6] 0.475 1.586 - -
MiniCPM-V-2_6 [75] 0.548 1.386 0.545 0.710
InternVL2-8B [10] 0.547 1.403 0.548 0.701
Ovis1.5-Gemma2-9B [50] 0.521 1.487 0.524 0.808
RBDash_72b [57] 0.525 1.413 0.527 0.740
MMAlaya2 [48] 0.537 1.423 0.541 0.747
Ovis1.5-Llama3-8B [50] 0.526 1.466 0.527 0.795
OmChat [82] 0.439 1.761 0.447 1.028
InternVL-Chat-V1-5 [11] 0.547 1.393 0.547 0.729
llava_onevision_qwen2_7b_si [37] 0.492 1.554 0.497 0.834
XComposer2d5 [11] 0.460 1.646 0.458 0.955
Pixtral-12B [3] 0.552 1.372 0.561 0.686
InternVL2-4B [10] 0.556 1.345 0.556 0.689
llava_onevision_qwen2_7b_ov [37] 0.488 1.590 0.490 0.861
glm-4v-9b [22] 0.511 1.463 0.527 0.730
molmo-7B-D-0924 [14] 0.519 1.490 0.523 0.813
XComposer2_4KHD [11] 0.433 1.781 0.435 1.075
MiniCPM-Llama3-V-2_5 [75] 0.526 1.432 0.530 0.767
VILA1.5-40b [40] 0.483 1.656 0.486 0.958
cambrian_34b [63] 0.546 1.377 0.549 0.701
WeMM [67] 0.488 1.565 0.479 0.888
Llama-3.2-11B-Vision-Instruct [52] 0.526 1.428 0.525 0.768
Qwen2-VL-2B-Instruct [65] 0.487 1.549 0.501 0.806
XComposer2 [10] 0.248 2.890 0.247 2.130
cogvlm2-llama3-chat-19B [26] 0.465 1.523 0.464 0.825
Idefics3-8B-Llama3 [35] 0.460 1.718 0.458 1.004
Mini-InternVL-Chat-4B-V1-5 [21] 0.527 1.433 0.525 0.765
XinYuan-VL-2B-Instruct [71] 0.395 1.951 0.392 1.240
llava_next_yi_34b [43] 0.476 1.636 0.483 0.883
InternVL2-2B [10] 0.476 1.563 0.483 0.848
Phi-3-Vision [1] 0.457 1.678 0.461 0.960
cambrian_13b [63] 0.509 1.526 0.506 0.852
Phi-3.5-Vision [1] 0.479 1.593 0.478 0.892
idefics2_8b [36] 0.449 1.761 0.453 1.015
cambrian_8b [63] 0.512 1.498 0.515 0.810
minimonkey [29] 0.450 1.673 0.455 0.939
Llama-3-MixSenseV1_1 [64] 0.474 1.615 0.477 0.922
llava_next_qwen_32b [43] 0.507 1.501 0.515 0.796
xgen-mm-phi3-interleave-r-v1.5 [73] 0.529 1.454 0.529 0.761
Eagle-X5-13B [59] 0.434 1.812 0.446 1.050
xgen-mm-phi3-dpo-r-v1.5 [73] 0.414 1.671 0.412 0.940
Mini-InternVL-Chat-2B-V1-5 [21] 0.428 1.709 0.440 0.973
llava_next_llama3 [43] 0.476 1.601 0.486 0.881
VILA1.5-13b [40] 0.485 1.695 0.480 1.006
Mantis-8B-siglip-llama3 [31] 0.411 1.833 0.409 1.116